1.) A sample has M = 70 and s = 4. In this sample, what is the X value corresponding to z = -1.00?

2.) Is it possible for a person A to have a larger X value than person B, but have a smaller z score?

3.) Consider the following set of data: [2, 2, 2, 2, 2, 2, 2, 2]. The standard deviation of this data will be:
Answer #1

1) In this sample,

The X value that corresponds to z = -1 is = M + z*s = 70 + (-1)*4 = 70 - 4 = 66

2) If

So, it is not possible.

3)Since all values are equal to 2, mean will also be equal to2

Since standard deviation = sqrt ( sum of squared deviation around mean) = sqrt(0) = 0